OpenStudy (anonymous):

How to write (x-1)(x+2)(x+3) +1 in standard form?

OpenStudy (anonymous):

wasn't this my example? :)

OpenStudy (anonymous):

yea..

OpenStudy (anonymous):

ok... let's do this with the first two binomials (x-1)(x+2). (x-1)(x+2) = x^2 - x + 2x - 2 = can you simplify the right side?

OpenStudy (anonymous):

i dont know how to

OpenStudy (anonymous):

combine like terms... like terms are the x's with the same exponent

OpenStudy (anonymous):

3^2 - 2 -2?

OpenStudy (anonymous):

 this is the what (x-1)(x+2) multiplied out